# Hessian stability and convergence rates for entropic and Sinkhorn potentials via semiconcavity

### v1

@@ 3.2. Quantitative stability estimates of Hessian
$\leq$ $\,\biggl{(}\frac{1}{\tau_{u}-\tau_{\ell}}+2\gamma_{\tau_{u}}\biggr{)}\,\frac{2\gamma_{\tau_{u}}}{\sqrt{2\pi}}\,\int_{0}^{\tau_{\ell}}\mathbb{E}[|Y^{\theta}_{s}|]\,\int_{s}^{\tau_{u}}\mathcal{I}^{\psi^{\nu}}(s,u)^{-\nicefrac{{1}}{{2}}}\mathrm{d}u\,\mathrm{d}s$
@@
By combining Section 3.2 with the above estimate and with Section 3.2 we finally deduce
$\mathbb{E}\|Z^{\theta}_{0}\|_{\mathrm{HS}}\leq\biggl{[}{\tau_{\ell}}^{-\nicefrac{{1}}{{2}}}+2\,{\tau_{\ell}}^{\nicefrac{{1}}{{2}}}\,\bar{\lambda}_{\tau_{\ell}}\biggr{]}\biggl{(}\int_{0}^{{\tau_{\ell}}}\mathbb{E}\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2}\mathrm{d}s\biggr{)}^{\nicefrac{{1}}{{2}}}+\int_{0}^{{\tau_{\ell}}}\mathbb{E}\|Z^{\theta}_{s}\|_{\mathrm{HS}}^{2}\mathrm{d}s$
$+\int_{0}^{{\tau_{\ell}}}\mathbb{E}\|\nabla^{3}\psi^{\nu}_{s}(X^{\psi^{\mu},\rho}_{s})\,Y^{\theta}_{s}\|_{\mathrm{HS}}\mathrm{d}s$
$\leq A\,\mathbf{W}_{2}(\mu,\nu)+\frac{K_{\delta^{\prime}\delta}^{\rho\nu}}{T^{2}}\,\mathbf{W}_{2}^{2}(\mu,\nu)\,,$
with
@@
Let us conclude by specifying Theorem 3.2 to diverse settings, relying on the explicit computations performed in Appendix A and Appendix B. Recall that hereafter we write $a\lesssim b$ whenever there exists a numerical constant $C>0$ (independent of $T,\nu,\rho,\mu$) such that $a\leq C\,b$.
[…]
Assume H 1, that $\mu\ll\nu$, and that both $\rho$ and $\nu$ are compactly supported in a ball of radius $R$ (big enough so that $R^{2}\geq T$). Then we have
Proof.
Since $\rho$ has compact support, the same computations performed in Section A.1 guarantee $\Lambda(\varphi^{\mu}_{0})<\infty$ and hence the validity of H 2. Our computations yield to
[…]
Assume H 1, that $\mu\ll\nu$, and that both $\rho$ and $\nu$ are log-concave, i.e., that their (negative) log-densities satisfy $\nabla^{2}U_{\rho}\geq\alpha_{\rho}$ and $\nabla^{2}U_{\nu}\geq\alpha_{\nu}$ for some $\alpha_{\rho},\,\alpha_{\nu}>0$ (w.l.o.g. such that $\alpha_{\rho}\vee\alpha_{\nu}<T^{-1}$). Then we have
Proof.
Since $\rho$ is log-concave, the same computations performed in Section A.2 guarantee $\Lambda(\varphi^{\mu}_{0})<\infty$ and hence the validity of H 2. Our computations yield to
Lastly, as we have already mentioned in the introduction, our explicit computations allow to straightforwardly deduce bounds akin Section 3.2 and Section 3.2 when considering one marginal log-concave and the other one with compact support.
3.3. Exponential convergence of Hessian of Sinkhorn iterates
We conclude with the proof of the convergence of gradient and Hessian of Sinkhorn iterates. This will be a straightforward application of our quantitative stability estimates and the explicit computations performed in Appendices A and B (this time considering the couple of marginals $(\rho,\mu)$ as fixed and considering $\mu^{n}$ as perturbation of $\mu$).
@@ 3.3. Exponential convergence of Hessian of Sinkhorn iterates
Under our assumptions, Talagrand inequality (TI($\tau$)) and the data processing inequality for relative entropy combined with [13, Theorem 1.2] and guarantee that
(3.6) $\mathbf{W}_{2}^{2}(\mu,\mu^{n+1,n})\leq 2\tau\,\mathscr{H}(\mu|\mu^{n+1,n})\leq 2\tau\,\mathscr{H}(\pi^{\mu}|\pi^{n+1,n})\leq 2\tau\bigg{(}1-\frac{T}{T+\tau\Lambda}\bigg{)}^{(n-N+1)}\mathscr{H}(\pi^{\mu}|\pi^{0,0})\,.$
In particular notice that this implies $\mathscr{H}(\mu|\mu^{n+1,n})<+\infty$, which combined with our assumption further guarantees the validity of H 2 for the marginals $\mu^{n+1,n}$ generated along Sinkhorn’s iterates. This allows us to apply Theorems 3.1 and 3.2 (with the couple $\nu,\mu$ there, replaced here as $\mu,\mu^{n+1,n}$) and deduce that
(3.7) $\|\nabla\varphi^{n+1}-\nabla\varphi^{\mu}\|^{2}_{\mathrm{L}^{2}(\rho)}\leq\frac{C_{\rho\mu}^{\delta}}{T^{2}}\,\mathbf{W}_{2}^{2}(\mu^{n+1,n},\mu)\,$
and
(3.8) $\|\nabla^{2}\varphi^{{n+1}}-\nabla^{2}\varphi^{\mu}\|_{\mathrm{L}^{1}(\rho)}\leq A\,\mathbf{W}_{2}(\mu^{n+1,n},\mu)+\frac{K_{\delta^{\prime}\delta}^{\rho\mu}}{T^{2}}\,\mathbf{W}_{2}^{2}(\mu^{n+1,n},\mu)\,,$
with $C_{\rho\mu}$, $A$ and $K_{\delta^{\prime}\delta}^{\rho\mu}$ defined as in the stability results, this time depending solely on $T$, and on the marginals $\rho$ and $\mu$. Putting together (3.6), (3.7) and (3.8) yields to
$\|\nabla\varphi^{n+1}-\nabla\varphi^{\mu}\|^{2}_{\mathrm{L}^{2}(\rho)}\lesssim\tau\bigg{(}1-\frac{T}{T+\tau\Lambda}\bigg{)}^{(n-N+1)}\mathscr{H}(\pi^{\mu}|\pi^{0,0})\,,$
$\|\nabla^{2}\varphi^{{n+1}}-\nabla^{2}\varphi^{\mu}\|_{\mathrm{L}^{1}(\rho)}\lesssim\sqrt{\tau}\bigg{(}1-\frac{T}{T+\tau\Lambda}\bigg{)}^{\frac{n-N+1}{2}}\,\sqrt{\mathscr{H}(\pi^{\mu}|\pi^{0,0})}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\tau\bigg{(}1-\frac{T}{T+\tau\Lambda}\bigg{)}^{(n-N+1)}\mathscr{H}(\pi^{\mu}|\pi^{0,0})\,.$
Finally, the specific asymptotics of the constants when $\rho$ and $\mu$ are compactly supported or log-concave can be obtained as already done in Sections 3.1, 3.1, 3.2 and 3.2 for the gradient and Hessian stability estimates. ∎
@@ Appendix A Explicit computations for $C^{\varphi^{\nu}}$ and $\Lambda(\varphi^{\nu}_{0})$
In this section we specify the constants appearing in the entropic stability bound of Theorem 2.1 to various settings. Before actually doing it, let us preliminary recall the well-known identities [21, 31, 15, 20]
@@
where $(X^{\psi^{\nu},\rho}_{s})_{s\in[0,T]}$ is the forward Schrödinger process (from $\rho$ to $\nu$) defined at (1.3), whereas $\mathrm{Cov}(X^{\psi^{\nu},\rho}_{0}|X^{\psi^{\nu},\rho}_{s}=y)$ is the covariance of the law of this process at initial time conditioned on being in $y$ at time $s$. This can be easily seen by recalling that
@@
and computing the Hessian as done in [15, Proposition 17] for the case $s=T$.
@@
where $(X^{\varphi^{\nu},\nu}_{s})_{s\in[0,T]}$ is the backward Schrödinger process defined at (1.4) and $\mathrm{Cov}(X^{\varphi^{\nu},\nu}_{0}|X^{\varphi^{\nu},\nu}_{s}=y)$ is the covariance of the law of this process at initial time conditioned on being in $y$ at time $s$.
Furthermore, let us recall here the following convexity backpropagation result along Hamilton-Jacobi-Bellman equations (see for instance [20, Lemma 3.1])
[…]
@@ A.1. Marginal $\rho$ with compact support
Next, let us compute $C^{\varphi^{\nu}}$ defined at (2.8). This can be easily accomplished since for any $l\leq u<T$ we have
@@ A.2. Log-concavity of $\rho$
Then, it is well known (cf. [15]) that $\nabla^{2}\varphi^{\nu}\geq\alpha_{\rho}-T^{-1}$ and hence we can take $\lambda(\varphi^{\nu}_{T})=\alpha_{\rho}-T^{-1}$. This is enough to deduce from Appendix A that
@@
We are now ready to compute $C^{\varphi^{\nu}}$ defined at (2.8). This can be easily accomplished since for any $l\leq u<T$ we have
@@
Appendix B Explicit computations for Hessians stability
In this section we will compute the constants appearing in our main general result Theorem 3.2 in two specific settings and analyze their behavior w.r.t. the parameters $T,\nu,\rho$ Hereafter we write $a\lesssim b$ whenever there exists a numerical constant $C>0$ (independent of $T,\nu,\rho$) such that $a\leq C\,b$. In order to compute the constants appearing in the Hessian stability bounds recall that $C^{\psi^{\nu}}_{\delta^{\prime},\delta}$ was introduced in (3.1) as
@@ B.1. Marginal $\nu$ with compact support
=(B.3)${=}\frac{2\,R^{2}}{T(1-\delta^{\prime})^{2}}\biggl{(}1-\exp\biggl{(}-\frac{R^{2}}{T}\,\frac{1}{\Lambda(\psi^{\nu}_{0})(1-\delta^{\prime})}\biggr{)}\biggr{)}^{-1}=\frac{2\,R^{2}}{T(1-\delta^{\prime})^{2}}\biggl{(}1-\exp\biggl{(}-\frac{1}{1-\delta^{\prime}}\biggr{)}\biggr{)}^{-1}$

### v2

@@ 3.1. Quantitative stability estimates of Hessian
By combining Section 3.1 with the above estimate and with Section 3.1 we finally get our thesis with
@@
Our estimates depend on the two free parameters $\delta^{\prime}<\delta\in[0,1]$. A priori, one could simply optimize over their choice; however, this optimization heavily depends on the semiconcavity parameters. In the Appendix, in Section B.2 and Section B.2, we specify Theorem 3.6 to diverse settings by fixing appropriate choices for $\delta$ and $\delta^{\prime}$, proving the specialized bounds stated in Theorem 1.1.
We conclude with the proof of the convergence of gradient and Hessian of Sinkhorn iterates. This will be a straightforward application of our quantitative stability estimates.
@@
Under our assumptions, Talagrand inequality (TI($\tau$)) and the data processing inequality for relative entropy combined with [CCGT24, Theorem 1.2] guarantee that
(3.6) $\mathbf{W}_{2}^{2}(\mu,\mu^{n+1,n})\leq 2\tau\,\mathscr{H}(\mu|\mu^{n+1,n})\leq 2\tau\,\mathscr{H}(\pi^{\mu}|\pi^{n+1,n})\leq 2\bigg(1-\frac{T}{T+\tau\Lambda}\bigg)^{(n-N+1)}\tau\mathscr{H}(\pi^{\mu}|\pi^{0,0})\,.$
In particular $\mathscr{H}(\mu|\mu^{n+1,n})<+\infty$, and hence the validity of H 2 for the marginal $\mu^{n+1,n}$. This allows us to apply Theorems 3.1 and 3.6 (with the pair $\nu,\mu$ there, replaced here as $\mu,\mu^{n+1,n}$) and deduce
(3.7) $\|\nabla\varphi^{n+1}-\nabla\varphi^{\mu}\|^{2}_{\mathrm{L}^{2}(\rho)}\leq\frac{C_{\rho\mu}^{\delta}}{T^{2}}\,\mathbf{W}_{2}^{2}(\mu^{n+1,n},\mu)\,$,
$\|\nabla^{2}\varphi^{{n+1}}-\nabla^{2}\varphi^{\mu}\|_{\mathrm{L}^{1}(\rho)}\leq A\,\mathbf{W}_{2}(\mu^{n+1,n},\mu)+\frac{K_{\delta^{\prime}\delta}^{\rho\mu}}{T^{2}}\,\mathbf{W}_{2}^{2}(\mu^{n+1,n},\mu)\,$,
with $C_{\rho\mu}$, $A$ and $K_{\delta^{\prime}\delta}^{\rho\mu}$ defined as in the stability results, this time depending solely on $T$, and on the marginals $\rho$ and $\mu$. Putting together (3.6) and (3.7) leads to our general bounds. Finally, the specific value of the uniform semiconcavity parameter $\Lambda$ and the asymptotics of the constants when $\rho$ and $\mu$ are compactly supported or log-concave can be obtained from the explicit computations performed in the Appendix (considering the pair of marginals $(\rho,\mu)$ as fixed and with $\mu^{n}$ seen as perturbation of $\mu$). ∎
@@ Appendix A Explicit computations for $C^{\varphi^{\nu}}$ and $\Lambda(\varphi^{\nu}_{0})$
In this section we specify the constants appearing in the entropic stability bound of Theorem 2.1 to various settings. Before actually doing it, let us preliminary recall the well-known identities [CP23, FGP20, CDG23, Con24]
@@
where $(X^{\psi^{\nu},\rho}_{s})_{s\in[0,T]}$ is the forward Schrödinger process (from $\rho$ to $\nu$) which we recall here to be defined as
$\mathrm{d}X^{\psi^{\nu},\rho}_{s}=-\nabla\psi^{\nu}_{s}(X^{\psi^{\nu},\rho}_{s})\mathrm{d}s+\mathrm{d}B_{s}\,,\quad X_{0}^{\psi^{\nu},\rho}\sim\rho\,$,
whereas $\mathrm{Cov}(X^{\psi^{\nu},\rho}_{0}|X^{\psi^{\nu},\rho}_{s}=y)$ is the covariance of the law of this process at initial time conditioned on being in $y$ at time $s$. This can be easily seen by recalling that
@@
and computing the Hessian as done in [CDG23, Proposition 17] for the case $s=T$.
@@
where $(X^{\varphi^{\nu},\nu}_{s})_{s\in[0,T]}$ is the backward Schrödinger process defined as
$\mathrm{d}X^{\varphi^{\nu},\nu}_{s}=-\nabla\varphi^{\nu}_{s}(X^{\varphi^{\nu},\nu}_{s})\mathrm{d}s+\mathrm{d}B_{s}\,,\quad X^{\varphi^{\nu},\nu}_{0}\sim\nu\,$,
and $\mathrm{Cov}(X^{\varphi^{\nu},\nu}_{0}|X^{\varphi^{\nu},\nu}_{s}=y)$ is the covariance of the law of this process at initial time conditioned on being in $y$ at time $s$.
Furthermore, let us recall here the following convexity backpropagation result along Hamilton-Jacobi-Bellman equations (see for instance [Con24, Lemma 3.1])
Lemma A.1.
@@ A.1. Marginal $\rho$ with compact support
Next, let us compute $C^{\varphi^{\nu}}$ defined as
$C^{\varphi^{\nu}}\coloneqq T\biggl(\int_{0}^{T}e^{\int_{0}^{s}2\lambda(\varphi^{\nu}_{t})\mathrm{d}t}\mathrm{d}s\biggr)^{-1}\,.$
This can be easily accomplished since for any $l\leq u<T$ we have
@@ A.2. Log-concavity of $\rho$
Then, it is well known [CDG23] that $\nabla^{2}\varphi^{\nu}\geq\alpha_{\rho}-T^{-1}$ and hence we can take $\lambda(\varphi^{\nu}_{T})=\alpha_{\rho}-T^{-1}$. This is enough to deduce from Appendix A that
@@
We are now ready to compute
$C^{\varphi^{\nu}}\coloneqq T\biggl(\int_{0}^{T}e^{\int_{0}^{s}2\lambda(\varphi^{\nu}_{t})\mathrm{d}t}\mathrm{d}s\biggr)^{-1}\,.$
This can be easily accomplished since for any $l\leq u<T$ we have
@@
Appendix B Explicit computations for the stability of Hessians
In this section we will compute the constants appearing in Theorem 3.1 and Theorem 3.6 in two specific settings and analyze their behavior w.r.t. the parameters $T,\nu,\rho$ . Hereafter we write $a\lesssim b$ whenever there exists a numerical constant $C>0$ (independent of $T,\nu,\rho$) such that $a\leq C\,b$. In order to compute the constants appearing in the stability bounds for the Hessian recall that $C^{\psi^{\nu}}_{\delta^{\prime},\delta}$ was introduced as
@@ B.1. Marginal $\nu$ with compact support
[…]
@@ B.2. Log-concavity of $\nu$
In the following table we summarize the values of the constants so far computed (up to numerical prefactors).
@@
Let us conclude by specifying our stability results of gradients and Hessians from Theorem 3.1 and Theorem 3.6 to the two settings considered above, relying on the explicit computations performed in Appendix A and Appendix B. This will prove the asymptotic bounds stated in Theorem 1.1 in the main article. Recall that hereafter we write $a\lesssim b$ whenever there exists a numerical constant $C>0$ (independent of $T,\nu,\rho,\mu$) such that $a\leq C\,b$.
Corollary B.1 (Stability of gradients for compactly supported marginals).
Assume H 1, that $\mathscr{H}(\nu|\mathrm{Leb})<\infty$, that both $\rho$ and $\nu$ are compactly supported in a ball of radius $R$ (big enough so that $R^{2}\geq T$) and that either $\mu\ll\nu$ or $\operatorname{supp}(\mu)\subseteq B_{R}(0)$. Then we have
Proof.
Firstly, assume $\mu\ll\nu$. Since $\rho$ has compact support, the same computations performed in Section A.1 guarantee $\Lambda(\varphi^{\mu}_{0})<\infty$ and hence the validity of the gradient estimates from Theorem 3.1. Our choice of $\delta$ in (B.2) and the following computations yield to $C_{\rho\nu}^{\delta}\lesssim R^{4}\,T^{-2}$.
Next, if we assume $\operatorname{supp}(\mu)\subseteq B_{R}(0)$ instead of $\mu\ll\nu$ we proceed as follows. Consider the probability measure $\nu^{\eta}\propto\mu+\eta\nu$ and observe that $\nu\ll\nu^{\eta}$ and $\mu\ll\nu^{\eta}$. The former implies the applicability of the previous case to the pair $(\nu^{\eta},\nu)$, which guarantees $\nabla\varphi^{\nu^{\eta}}\rightarrow\nabla\varphi^{\nu}$ in $\mathrm{L}^{2}(\rho)$ as the regularization parameter $\eta$ vanishes. The latter ensures the validity of quantitative stability bound for the pair $(\mu,\nu^{\eta})$, uniformly in $\eta>0$, since $\operatorname{supp}(\nu^{\eta})\subseteq B_{R}(0)$ for all $\eta\geq 0$. Thus, it is sufficient to apply the triangle inequality and eventually send $\eta\downarrow 0$ to recover the final stability bound for the original pair $(\mu,\nu)$. ∎
Corollary B.2 (Stability of gradients for log-concave marginals).
Assume H 1, $\mathscr{H}(\nu|\mathrm{Leb})<\infty$ and that both $\rho$ and $\nu$ are log-concave, i.e., that their (negative) log-densities satisfy $\nabla^{2}U_{\rho}\geq\alpha_{\rho}$ and $\nabla^{2}U_{\nu}\geq\alpha_{\nu}$ for some $\alpha_{\rho},\,\alpha_{\nu}>0$ (w.l.o.g. such that $\alpha_{\rho}\vee\alpha_{\nu}<T^{-1}$). Then we have
Proof.
Firstly, let us consider the case where $\mu\ll\nu$. Since $\rho$ has log-concave density, the same computations performed in Section A.2 guarantee $\Lambda(\varphi^{\mu}_{0})<\infty$ and hence the validity of the validity of the gradient estimates from Theorem 3.1. Our choice of $\delta$ in (B.2) and the following computations yield to $C_{\rho\nu}^{\delta}\lesssim\alpha_{\rho}^{-1}\,\alpha_{\nu}^{-1}\,T^{-2}$.
To lift the assumption $\mu\ll\nu$ we argue in the following way. Fix a regularization parameter $\eta>0$ and consider the convolution with the heat kernel $\nu^{\eta}$. Then, as the (negative) log-density of $\nu^{\eta}$ satisfies $\nabla^{2}U_{\nu^{\eta}}\geq\frac{\alpha_{\nu}}{1+\eta\alpha_{\nu}}$ provided $\nabla^{2}U_{\nu}\geq\alpha_{\nu}$ (see for instance [HÅ06, Theorem 1]), the log-concavity parameter of $\nu^{\eta}$ converges to the one of $\nu$ as $\eta\downarrow 0$. Since $\nu\ll\nu^{\eta}$, we can thus apply the previous result and deduce that as $\eta\downarrow 0$ we have $\nabla\varphi^{\nu^{\eta}}\rightarrow\nabla\varphi^{\nu}$ in $\mathrm{L}^{2}(\rho)$. Moreover, since $\nu^{\eta}\sim\mathrm{Leb}$ and $\mu\ll\mathrm{Leb}$, we can again apply the previous result to the pair $(\mu,\nu^{\eta})$. Then, it is enough sending $\eta\downarrow 0$ to recover the final stability bound for the original pair $(\mu,\nu)$. ∎
Similarly, for the Hessian bounds we have the followings.
Corollary B.3 (Stability of Hessians for compactly supported marginals).
Assume H 1, that $\mathscr{H}(\nu|\mathrm{Leb})<\infty$, that both $\rho$ and $\nu$ are compactly supported in a ball of radius $R$ (big enough so that $R^{2}\geq T$) and that either $\mu\ll\nu$ or $\operatorname{supp}(\mu)\subseteq B_{R}(0)$. Then we have
Proof.
Firstly, we prove this result in the case $\mu\ll\nu$. Since $\rho$ has compact support, the same computations performed in Section A.1 guarantee $\Lambda(\varphi^{\mu}_{0})<\infty$ and hence the validity of of our general stability estimates for the Hessian of Schrödinger potentials from Theorem 3.6. Our computations yield to
The proof of this result when considering the assumption $\operatorname{supp}(\mu)\subseteq B_{R}(0)$ (instead of $\mu\ll\nu$) can be obtained via the same regularization procedure described in the proof of Section B.2 and for this reason we omit it here. ∎
Corollary B.4 (Stability of Hessians for log-concave marginals).
Assume H 1, $\mathscr{H}(\nu|\mathrm{Leb})<\infty$ and that both $\rho$ and $\nu$ are log-concave, i.e., that their (negative) log-densities satisfy $\nabla^{2}U_{\rho}\geq\alpha_{\rho}$ and $\nabla^{2}U_{\nu}\geq\alpha_{\nu}$ for some $\alpha_{\rho},\,\alpha_{\nu}>0$ (w.l.o.g. such that $\alpha_{\rho}\vee\alpha_{\nu}<T^{-1}$). Then we have
Proof.
We prove this result under the additional assumption $\mu\ll\nu$. The general result can be obtained following the same regularization procedure considered in the proof of Section B.2. Since $\rho$ is log-concave, the same computations performed in Section A.2 guarantee $\Lambda(\varphi^{\mu}_{0})<\infty$ and hence the validity of our general stability estimates for the Hessian of Schrödinger potentials from Theorem 3.6. Our computations yield to